# An optimal parallel-in-time preconditioner for parabolic optimal control problems

### v1

@@ 2 Main Result
$\begin{bmatrix}I_{mn}&-\Sigma^{*}\\ \Sigma&I_{mn}\end{bmatrix}$ $=$ $\mathcal{Q}\begin{bmatrix}I_{mn}+\mathbf{i}\Sigma&\\ &I_{mn}-\mathbf{i}\Sigma\end{bmatrix}\mathcal{Q}^{*}.$
@@
~$\widetilde{\mathcal{H}}^{-1}\widetilde{\mathcal{B}}$ $=$ $\begin{bmatrix}\mathcal{V}&\\ &\mathcal{U}\end{bmatrix}\mathcal{Q}\begin{bmatrix}I_{mn}+\mathbf{i}\Sigma&\\ &I_{mn}-\mathbf{i}\Sigma\end{bmatrix}\mathcal{Q}^{*}\begin{bmatrix}\mathcal{V}^{*}&\\ &\mathcal{U}^{*}\end{bmatrix}\mathcal{Q}$
@@
Clearly, the eigenvalues of ~$\widetilde{\mathcal{H}}^{-1}\widetilde{\mathcal{B}}$ locates in $\{1+{\bf i}x|x\in[-||\widetilde{\mathcal{E}}||_{2},||\widetilde{\mathcal{E}}||_{2}]\}\subset\mathbb{C}$. As the matrix $\widetilde{\mathcal{T}}$ is positive definite (i.e., $\widetilde{\mathcal{T}}+\widetilde{\mathcal{T}}^{\top}$ is Hermitian positive definite), we know from [3] that ~$\|\widetilde{\mathcal{E}}\|_{2}<1$. Therefore,
@@
Note that $\widetilde{\mathcal{C}}_{\epsilon}+\alpha{I}_{mn}=\widetilde{\mathcal{T}}+\alpha{I}_{mn}-\epsilon E_{1}E_{n}^{\top}$, where $E_{i}=e_{i}\otimes I_{m}$ with $e_{i}$ denoting the $i$-th column of $I_{n}$ and $\otimes$ denoting the Kronecker product. By the Sherman–Morrison–Woodbury formula, we have
@@
$(\widetilde{\mathcal{T}}^{\top}+\alpha{I}_{mn})^{-1}E_{n}\mathcal{Z}_{\epsilon}^{-1}E_{1}^{\top}$ $=\begin{bmatrix}T_{o}^{-n}\mathcal{Z}_{\epsilon}^{-1}&&&\\ T_{o}^{-n+1}\mathcal{Z}_{\epsilon}^{-1}&&&\\ \vdots&&&\\ T_{o}^{-1}\mathcal{Z}_{\epsilon}^{-1}&&&\end{bmatrix},$
(22) $(\widetilde{\mathcal{T}}+\alpha{I}_{mn})^{-1}E_{1}\mathcal{Z}_{\epsilon}^{-1}E_{n}^{\top}$ $=\begin{bmatrix}&&&T_{0}^{-1}\mathcal{Z}_{\epsilon}^{-1}\\ &&&T_{0}^{-2}\mathcal{Z}_{\epsilon}^{-1}\\ &&&\vdots\\ &&&T_{0}^{-n}\mathcal{Z}_{\epsilon}^{-1}\end{bmatrix}.$
@@
$=\begin{bmatrix}&&&T_{0}^{-1}\mathcal{Z}_{\epsilon}^{-1}\\ &&&T_{0}^{-2}\mathcal{Z}_{\epsilon}^{-1}\\ &&&\vdots\\ &&&T_{0}^{-n}\mathcal{Z}_{\epsilon}^{-1}\end{bmatrix}.$
@@
$=$ $\epsilon(I_{n}\otimes\mathcal{Q}_{m})\begin{bmatrix}&&&\Lambda^{1}(I_{m}-\epsilon\Lambda^{n})^{-1}\\ &&&\Lambda^{2}(I_{m}-\epsilon\Lambda^{n})^{-1}\\ &&&\vdots\\ &&&\Lambda^{n}(I_{m}-\epsilon\Lambda^{n})^{-1}\end{bmatrix}(I_{n}\otimes\mathcal{Q}_{m}^{\top}).$
@@
$\leq$ $\epsilon\sqrt{\left\|\sum_{j=1}^{n}(\Lambda^{j}(I_{m}-\epsilon\Lambda^{n})^{-1})^{2}\right\|_{2}}$
$=$ $\epsilon\sqrt{\max_{1\leq k\leq m}\sum_{j=1}^{n}\left(\frac{\lambda_{k}^{j}}{1-\epsilon\lambda_{k}^{n}}\right)^{2}}\leq\epsilon\sqrt{\sup_{\lambda\in(0,1)}\sum_{j=1}^{n}\left(\frac{\lambda^{j}}{1-\epsilon\lambda^{n}}\right)^{2}},$
where the last inequality comes from the fact that $\sigma(T_{0}^{-1})\subset(0,1)$. Since the functions $g_{j}(x)=\frac{x^{j}}{1-\epsilon x^{n}}$ are monotonically increasing on $x\in[0,1]$ for each $j=1,2,...,n$, we have
$\|(\widetilde{\mathcal{C}}_{\epsilon}+\alpha{I}_{mn})^{-1}(\widetilde{\mathcal{T}}+\alpha{I}_{mn})-I_{mn}\|_{2}$$\leq$ $\epsilon\sqrt{\sum_{j=1}^{n}\frac{1}{(1-\epsilon)^{2}}}$
@@
Note that the result and the proof of $\|(\widetilde{\mathcal{C}}_{\epsilon}^{\top}+\alpha{I}_{mn})^{-1}(\widetilde{\mathcal{T}}^{\top}+\alpha{I}_{mn})\|_{2}$ is the same as Lemma 2.11. Given any constant $\eta\in(0,1)$, for any $\epsilon\in(0,\eta]$, we have
@@
$=$ $\frac{1}{2}\begin{bmatrix}\mathcal{R}_{1}+\mathcal{R}_{2}^{\top}&(\mathcal{R}_{2}\widetilde{\mathcal{E}}-\widetilde{\mathcal{E}}\mathcal{R}_{1}^{\top})^{\top}\\ \mathcal{R}_{2}\widetilde{\mathcal{E}}-\widetilde{\mathcal{E}}\mathcal{R}_{1}^{\top}&\mathcal{R}_{2}+\mathcal{R}_{2}^{\top}\end{bmatrix}$
$=$ $\frac{1}{2}\begin{bmatrix}\mathcal{R}_{1}+\mathcal{R}_{2}^{\top}&\\ &\mathcal{R}_{2}+\mathcal{R}_{2}^{\top}\end{bmatrix}+\frac{1}{2}\begin{bmatrix}&(\mathcal{R}_{2}\widetilde{\mathcal{E}}-\widetilde{\mathcal{E}}\mathcal{R}_{1}^{\top})^{\top}\\ \mathcal{R}_{2}\widetilde{\mathcal{E}}-\widetilde{\mathcal{E}}\mathcal{R}_{1}^{\top}&\end{bmatrix}.$
@@
$\frac{1}{2}\left\|\begin{bmatrix}&(\mathcal{R}_{2}\mathcal{E}-\mathcal{E}\mathcal{R}_{1}^{\top})^{\top}\\ \mathcal{R}_{2}\mathcal{E}-\mathcal{E}\mathcal{R}_{1}^{\top}&\end{bmatrix}\right\|_{2}$
$\leq$ $\frac{1}{2}\max\{\|(\mathcal{R}_{2}\mathcal{E}-\mathcal{E}\mathcal{R}_{1}^{\top})^{\top}\|_{2},\|\mathcal{R}_{2}\mathcal{E}-\mathcal{E}\mathcal{R}_{1}^{\top}\|_{2}\}$
$=$ $\frac{1}{2}\|\mathcal{R}_{2}\mathcal{E}-\mathcal{E}\mathcal{R}_{1}^{\top}\|_{2}$
$\leq$ $\frac{1}{2}(\|\mathcal{R}_{2}\|_{2}\|\mathcal{E}\|_{2}+\|\mathcal{E}\|_{2}\|\mathcal{R}_{1}\|_{2})$
@@
Given $\delta\in(0,1)$, for any $\epsilon\in(0,c_{\tau}]$, where $c_{\tau}:=\frac{\delta\sqrt{\tau}}{\delta\sqrt{\tau}+2\sqrt{T}}$, the residuals of the iterates generated by applying GMRES to solving the auxiliary linear system (16) satisfy
@@
(~$\mathbb{H}(\widetilde{\mathcal{H}}_{\epsilon}^{-1}\widetilde{\mathcal{B}})$ $=$ $I_{2mn}+\mathcal{H}(\widetilde{\mathcal{H}}_{\epsilon}^{-1}\widetilde{\mathcal{B}}-I_{2mn})$
$\succeq$ $(1-\delta)I_{2mn}$
@@
Given $\delta\in(0,1)$, for any $\epsilon\in(0,c_{\tau}]$, where $c_{\tau}:=\frac{\delta\sqrt{\tau}}{\delta\sqrt{\tau}+2\sqrt{T}}$, the residuals of the iterates generated by applying GMRES to solving the preconditioned linear system (15) satisfy
@@
Let $\tilde{\mathbf{x}}_{0}=\mathcal{W}^{\frac{1}{2}}\mathcal{G}\mathbf{x}_{0}$. Then, it holds that $\mathbf{x}_{0}=\mathcal{G}^{-1}\mathcal{W}^{-\frac{1}{2}}\tilde{\mathbf{x}}_{0}$. Take $\tilde{\mathbf{x}}_{0}$ as an initial guess of GMRES solver for solving the auxiliary linear system (16). Then, according to Theorem 2.2, we see that
$\left\|\mathbf{r}_{k}\right\|_{2}\leq\sqrt{2}||\mathcal{W}^{-\frac{1}{2}}||_{2}||\tilde{\bf r}_{k}||_{2},\quad k=1,2,...,$
where $\tilde{\mathbf{r}}_{k}:=\tilde{\bf b}-\widetilde{\mathcal{H}}_{\epsilon}^{-1}\widetilde{\mathcal{B}}\tilde{\mathbf{x}}_{k}$ denotes the residual vector at the $k$-th GMRES iteration for solving (16) and $\tilde{\mathbf{x}}_{k}$ denotes the $k$-th iterative solution of GMRES solver for solving (16). According to Lemma 2.19, we can further estimate $||\tilde{\bf r}_{k}||_{2}$ as
@@
As a consequence of Theorem 2.21, GMRES can achieve a mesh-independent convergence rate when $\epsilon=\mathcal{O}(\tau^{1/2})$.
@@ 2.1 Implementation
First, we discuss the computation of $\mathcal{A}\mathbf{v}$ for any given vector $\mathbf{v}$. The computation of matrix-vector product $\mathcal{A}\mathbf{v}$ can be computed in $\mathcal{O}(mn)$ since $\mathcal{A}$ is a sparse matrix consisting of two simple bi-diagonal block Toeplitz matrices. The required storage is of $\mathcal{O}(mn)$.
At each GMRES iteration, the matrix-vector product $\mathcal{P}_{\epsilon}^{-1}\mathbf{v}$ for a given vector $\mathbf{v}$ needs to be computed. Recalling that $\epsilon$-circulant matrices are diagonalizable by the product of a diagonal matrix and a discrete Fourier matrix $\mathbb{F}_{n}=\frac{1}{\sqrt{n}}[\theta_{n}^{(i-1)(j-1)}]_{i,j=1}^{n}\in\mathbb{C}^{n\times n}$ with $\theta_{n}=\exp{(\frac{2\pi\mathbf{i}}{n})}$, we can represent the matrix $C_{\epsilon,n}$ defined by (13) using the diagonalization $C_{\epsilon,n}=D_{\epsilon}^{-1}\mathbb{F}_{n}\Lambda_{\epsilon,n}\mathbb{F}_{n}^{*}D_{\epsilon}$. Note that $\Lambda_{\epsilon,n}$ is a diagonal matrix.
@@
$=$ $\mathcal{U}\begin{bmatrix}(\Lambda^{*}_{\epsilon,n}+\alpha I_{n})\otimes M_{m}+\tau I_{n}\otimes(-L_{m})&\\ &(\Lambda_{\epsilon,n}+\alpha I_{n})\otimes M_{m}+\tau I_{n}\otimes(-L_{m})\end{bmatrix}\mathcal{U}^{-1}$
@@
Note that $\mathcal{U}=\begin{bmatrix}(\mathbb{F}_{n}^{*}D_{\epsilon})^{\top}\otimes I_{m}&\\ &D_{\epsilon}^{-1}\mathbb{F}_{n}\otimes I_{m}\end{bmatrix}$, where $D_{\epsilon}={\rm diag}(\epsilon^{\frac{i-1}{N}})_{i=1}^{n}$ and $\mathbb{F}_{n}=\frac{1}{\sqrt{n}}[\theta_{n}^{(i-1)(j-1)}]_{i,j=1}^{n}$ with $\theta_{n}=\exp(\frac{2\pi{\bf i}}{n})$. Also, $\Lambda_{\epsilon,n}={\rm diag}(\lambda_{i-1}^{(\epsilon)})_{i=1}^{n}$ with $\lambda_{k}^{(\epsilon)}=1-\epsilon^{\frac{1}{n}}\theta_{n}^{-k}$.
@@
    $\widetilde{\mathbf{v}}=\begin{bmatrix}(\Lambda^{*}_{\epsilon,n}+\alpha I_{n})\otimes M_{m}+\tau I_{n}\otimes(-L_{m})&\\ &(\Lambda_{\epsilon,n}+\alpha I_{n})\otimes M_{m}+\tau I_{n}\otimes(-L_{m})\end{bmatrix}^{-1}\widehat{\mathbf{v}},$

### v2

@@ 2 Main Result
$\begin{bmatrix}I_{mn}&-{\color[rgb]{0,0,0}\Sigma}\\ \Sigma&I_{mn}\end{bmatrix}$ $=$ $\mathcal{Q}\begin{bmatrix}I_{mn}+\mathbf{i}\Sigma&\\ &I_{mn}-\mathbf{i}\Sigma\end{bmatrix}\mathcal{Q}^{*}.$
@@
~$\widetilde{\mathcal{H}}^{-1}\widetilde{\mathcal{B}}$ $=$ $\begin{bmatrix}\mathcal{V}&\\ &\mathcal{U}\end{bmatrix}\mathcal{Q}\begin{bmatrix}I_{mn}+\mathbf{i}\Sigma&\\ &I_{mn}-\mathbf{i}\Sigma\end{bmatrix}\mathcal{Q}^{*}{\color[rgb]{0,0,0}\begin{bmatrix}\mathcal{V}^{{\color[rgb]{0,0,0}\top}}&\\ &\mathcal{U}^{{\color[rgb]{0,0,0}\top}}\end{bmatrix}}$
@@
Clearly, the eigenvalues of ~$\widetilde{\mathcal{H}}^{-1}\widetilde{\mathcal{B}}$ locates in $\{1+{\bf i}x|x\in[-||\widetilde{\mathcal{E}}||_{2},||\widetilde{\mathcal{E}}||_{2}]\}\subset\mathbb{C}$. As the matrix $\widetilde{\mathcal{T}}$ is positive definite (i.e., $\widetilde{\mathcal{T}}+\widetilde{\mathcal{T}}^{\top}$ is symmetric positive definite), we know from [3, Theorem 2.2] that ~$\|\widetilde{\mathcal{E}}\|_{2}<1$. Namely, ~$\widetilde{\mathcal{E}}=:V(\widetilde{\mathcal{T}},\alpha)$, which is defined as the so-called extrapolated Cayley transform in [3], is strictly less than one in the Euclidean norm. For more details, refer to the proof of [3, Theorem 2.2]. Therefore,
@@
$(\widetilde{\mathcal{C}}_{\epsilon}+\alpha{I}_{mn})^{-1}(\widetilde{\mathcal{T}}+\alpha{I}_{mn})=I_{mn}+(\widetilde{\mathcal{T}}+\alpha{I}_{mn})^{-1}E_{1}\mathcal{Z}_{\epsilon}^{-1}E_{n}^{\top},$
where $E_{i}=e_{i}\otimes I_{m}$ with $e_{i}$ denoting the $i$-th column of $I_{n}$ and $\otimes$ denoting the Kronecker product.
@@
Note that $\widetilde{\mathcal{C}}_{\epsilon}+\alpha{I}_{mn}=\widetilde{\mathcal{T}}+\alpha{I}_{mn}-\epsilon E_{1}E_{n}^{\top}$. By the Sherman–Morrison–Woodbury formula, we have
@@
$(\widetilde{\mathcal{T}}^{\top}+\alpha{I}_{mn})^{-1}E_{n}\mathcal{Z}_{\epsilon}^{-1}E_{1}^{\top}$ $={\color[rgb]{0,0,0}\begin{bmatrix}T_{o}^{-n}\mathcal{Z}_{\epsilon}^{-1}&0&\cdots&0\\ T_{o}^{-n+1}\mathcal{Z}_{\epsilon}^{-1}&\vdots&&\vdots\\ \vdots&\vdots&&\vdots\\ T_{o}^{-1}\mathcal{Z}_{\epsilon}^{-1}&0&\cdots&0\end{bmatrix}},$
(23) $(\widetilde{\mathcal{T}}+\alpha{I}_{mn})^{-1}E_{1}\mathcal{Z}_{\epsilon}^{-1}E_{n}^{\top}$ $={\color[rgb]{0,0,0}\begin{bmatrix}0&\cdots&0&T_{0}^{-1}\mathcal{Z}_{\epsilon}^{-1}\\ \vdots&&\vdots&T_{0}^{-2}\mathcal{Z}_{\epsilon}^{-1}\\ \vdots&&\vdots&\vdots\\ 0&\cdots&0&T_{0}^{-n}\mathcal{Z}_{\epsilon}^{-1}\end{bmatrix}}.$
@@
$={\color[rgb]{0,0,0}\begin{bmatrix}0&\cdots&0&T_{0}^{-1}\mathcal{Z}_{\epsilon}^{-1}\\ \vdots&&\vdots&T_{0}^{-2}\mathcal{Z}_{\epsilon}^{-1}\\ \vdots&&\vdots&\vdots\\ 0&\cdots&0&T_{0}^{-n}\mathcal{Z}_{\epsilon}^{-1}\end{bmatrix}}.$
@@
$=$ $\epsilon(I_{n}\otimes\mathcal{Q}_{m}){\color[rgb]{0,0,0}\begin{bmatrix}0&\cdots&0&\Lambda^{1}(I_{m}-\epsilon\Lambda^{n})^{-1}\\ \vdots&&\vdots&\Lambda^{2}(I_{m}-\epsilon\Lambda^{n})^{-1}\\ \vdots&&\vdots&\vdots\\ 0&\cdots&0&\Lambda^{n}(I_{m}-\epsilon\Lambda^{n})^{-1}\end{bmatrix}}(I_{n}\otimes\mathcal{Q}_{m}^{\top}).$
@@
$\leq$ $\epsilon{\color[rgb]{0,0,0}\cdot}\sqrt{\left\|\sum_{j=1}^{n}(\Lambda^{j}(I_{m}-\epsilon\Lambda^{n})^{-1})^{2}\right\|_{2}}$
$=$ $\epsilon{\color[rgb]{0,0,0}\cdot}\sqrt{\max_{1\leq k\leq m}\sum_{j=1}^{n}\left(\frac{\lambda_{k}^{j}}{1-\epsilon\lambda_{k}^{n}}\right)^{2}}\leq\epsilon{\color[rgb]{0,0,0}\cdot}\sqrt{\sup_{\lambda\in(0,1)}\sum_{j=1}^{n}\left(\frac{\lambda^{j}}{1-\epsilon\lambda^{n}}\right)^{2}},$
where the last inequality comes from the fact that $\sigma(T_{0}^{-1})\subset(0,1)$. Since the functions $g_{j}(x)=\frac{x^{j}}{1-\epsilon x^{n}}$ are monotonically increasing on $x\in[0,1]$ for each $j=1,2,\dots,n$, we have
$\|(\widetilde{\mathcal{C}}_{\epsilon}+\alpha{I}_{mn})^{-1}(\widetilde{\mathcal{T}}+\alpha{I}_{mn})-I_{mn}\|_{2}$$\leq$ $\epsilon{\color[rgb]{0,0,0}\cdot}\sqrt{\sum_{j=1}^{n}\frac{1}{(1-\epsilon)^{2}}}$
@@
Following on the proof presented in Lemma 2.11, we can derive the same bound for $\|(\widetilde{\mathcal{C}}_{\epsilon}^{\top}+\alpha{I}_{mn})^{-1}(\widetilde{\mathcal{T}}^{\top}+\alpha{I}_{mn})-I_{mn}\|_{2}$. Specifically, given any constant $\eta\in(0,1)$, for any $\epsilon\in(0,\eta]$, we have
@@
$=$ $\frac{1}{2}\begin{bmatrix}\mathcal{R}_{1}+{\color[rgb]{0,0,0}\mathcal{R}_{1}^{\top}}&(\mathcal{R}_{2}\widetilde{\mathcal{E}}-\widetilde{\mathcal{E}}\mathcal{R}_{1}^{\top})^{\top}\\ \mathcal{R}_{2}\widetilde{\mathcal{E}}-\widetilde{\mathcal{E}}\mathcal{R}_{1}^{\top}&\mathcal{R}_{2}+\mathcal{R}_{2}^{\top}\end{bmatrix}$
$=$ $\frac{1}{2}\begin{bmatrix}\mathcal{R}_{1}+{\color[rgb]{0,0,0}\mathcal{R}_{1}^{\top}}&\\ &\mathcal{R}_{2}+\mathcal{R}_{2}^{\top}\end{bmatrix}+\frac{1}{2}\begin{bmatrix}&(\mathcal{R}_{2}\widetilde{\mathcal{E}}-\widetilde{\mathcal{E}}\mathcal{R}_{1}^{\top})^{\top}\\ \mathcal{R}_{2}\widetilde{\mathcal{E}}-\widetilde{\mathcal{E}}\mathcal{R}_{1}^{\top}&\end{bmatrix}.$
@@
$\frac{1}{2}\left\|\begin{bmatrix}&(\mathcal{R}_{2}\widetilde{\mathcal{E}}-\widetilde{\mathcal{E}}\mathcal{R}_{1}^{\top})^{\top}\\ \mathcal{R}_{2}\widetilde{\mathcal{E}}-\widetilde{\mathcal{E}}\mathcal{R}_{1}^{\top}&\end{bmatrix}\right\|_{2}$
$=$ $\frac{1}{2}\max\{\|(\mathcal{R}_{2}\widetilde{\mathcal{E}}-\widetilde{\mathcal{E}}\mathcal{R}_{1}^{\top})^{\top}\|_{2},\|\mathcal{R}_{2}\widetilde{\mathcal{E}}-\widetilde{\mathcal{E}}\mathcal{R}_{1}^{\top}\|_{2}\}$
$=$ $\frac{1}{2}\|\mathcal{R}_{2}\widetilde{\mathcal{E}}-\widetilde{\mathcal{E}}\mathcal{R}_{1}^{\top}\|_{2}$
@@
Given $\delta\in(0,1)$, for any $\epsilon\in(0,c_{\tau}]$, where $c_{\tau}:=\frac{\delta\sqrt{\tau}}{\delta\sqrt{\tau}+2\sqrt{T}}$ with $\tau=T/(n+1)$, the residuals of the iterates generated by applying GMRES to solving the auxiliary linear system (17) satisfy
@@
(~$\mathbb{H}(\widetilde{\mathcal{H}}_{\epsilon}^{-1}\widetilde{\mathcal{B}})$ $=$ $I_{2mn}+{\color[rgb]{0,0,0}\mathbb{H}}(\widetilde{\mathcal{H}}_{\epsilon}^{-1}\widetilde{\mathcal{B}}-I_{2mn})$
$\succeq$ ${\color[rgb]{0,0,0}I_{2mn}-\|\mathbb{H}(\widetilde{\mathcal{H}}_{\epsilon}^{-1}\widetilde{\mathcal{B}}-I_{2mn})\|_{2}I_{2mn}}$
$=$ $(1-\delta)I_{2mn}$
@@
Given $\delta\in(0,1)$, for any $\epsilon\in(0,c_{\tau}]$, where $c_{\tau}:=\frac{\delta\sqrt{\tau}}{\delta\sqrt{\tau}+2\sqrt{T}}$ with $\tau=T/(n+1)$, the residuals of the iterates generated by applying GMRES to solving the preconditioned linear system (16) satisfy
@@
Let $\tilde{\mathbf{x}}_{0}=\mathcal{W}^{\frac{1}{2}}\mathcal{G}\mathbf{x}_{0}$. Then, it holds that $\mathbf{x}_{0}=\mathcal{G}^{-1}\mathcal{W}^{-\frac{1}{2}}\tilde{\mathbf{x}}_{0}$. Take $\tilde{\mathbf{x}}_{0}$ as an initial guess of the GMRES algorithm for solving the auxiliary linear system (17). Then, according to Theorem 2.2, we see that
$\left\|\mathbf{r}_{k}\right\|_{2}\leq\sqrt{2}||\mathcal{W}^{-\frac{1}{2}}||_{2}||\tilde{\bf r}_{k}||_{2},\quad k=1,2,\dots,$
where $\tilde{\mathbf{r}}_{k}:=\tilde{\bf b}-\widetilde{\mathcal{H}}_{\epsilon}^{-1}\widetilde{\mathcal{B}}\tilde{\mathbf{x}}_{k}$ denotes the residual vector at the $k$-th GMRES iteration for solving (17) and $\tilde{\mathbf{x}}_{k}$ denotes the $k$-th iterative solution of GMRES for solving (17). According to Lemma 2.19, we can further estimate $||\tilde{\bf r}_{k}||_{2}$ as
@@
As a consequence of Theorem 2.21, GMRES can achieve a mesh-independent convergence rate when $\epsilon=\mathcal{O}(\tau^{1/2})$. This is because the factor $\frac{\sqrt{-\delta^{2}+8\delta+2}}{2+\delta}$ lies in the interval $(0,1)$ for this choice of $\epsilon$. As the iteration count $k$ increases, $\|\mathbf{r}_{k}\|_{2}$ tends to zero uniformly.
@@ 2.1 Implementation
First, we discuss the computation of $\mathcal{A}\mathbf{v}$ for any given vector $\mathbf{v}$. The computation of the matrix-vector product $\mathcal{A}\mathbf{v}$ can be computed in $\mathcal{O}(mn)$ since $\mathcal{A}$ is a sparse matrix consisting of two simple bi-diagonal block Toeplitz matrices. The required storage is of $\mathcal{O}(mn)$.
At each GMRES iteration, the matrix-vector product $\mathcal{P}_{\epsilon}^{-1}\mathbf{v}$ for a given vector $\mathbf{v}$ needs to be computed. Recalling that $\epsilon$-circulant matrices are diagonalizable by the product of a diagonal matrix and a discrete Fourier matrix $\mathbb{F}_{n}=\frac{1}{\sqrt{n}}[\theta_{n}^{(i-1)(j-1)}]_{i,j=1}^{n}\in\mathbb{C}^{n\times n}$ with $\theta_{n}=\exp{(\frac{2\pi\mathbf{i}}{n})}$, we can represent the matrix $C_{\epsilon,n}$ defined by (13) using the diagonalization $C_{\epsilon,n}=D_{\epsilon}^{-1}\mathbb{F}_{n}\Lambda_{\epsilon,n}\mathbb{F}_{n}^{*}D_{\epsilon}$ , where $D_{\epsilon}={\rm diag}(\epsilon^{\frac{i-1}{N}})_{i=1}^{n}$. Note that $\Lambda_{\epsilon,n}$ is a diagonal matrix.
@@
$=$ $\mathcal{U}\begin{bmatrix}(\Lambda^{*}_{\epsilon,n}+\alpha I_{n})\otimes M_{m}+\tau I_{n}\otimes{\color[rgb]{0,0,0}K_{m}}&\\ &(\Lambda_{\epsilon,n}+\alpha I_{n})\otimes M_{m}+\tau I_{n}\otimes{\color[rgb]{0,0,0}K_{m}}\end{bmatrix}\mathcal{U}^{-1}$
@@
Note that $\mathcal{U}=\begin{bmatrix}(\mathbb{F}_{n}^{*}D_{\epsilon})^{\top}\otimes I_{m}&\\ &D_{\epsilon}^{-1}\mathbb{F}_{n}\otimes I_{m}\end{bmatrix}$ and $\Lambda_{\epsilon,n}={\rm diag}(\lambda_{i-1}^{(\epsilon)})_{i=1}^{n}$ with $\lambda_{k}^{(\epsilon)}=1-\epsilon^{\frac{1}{n}}\theta_{n}^{-k}$.
@@
    $\widetilde{\mathbf{v}}=\begin{bmatrix}(\Lambda^{*}_{\epsilon,n}+\alpha I_{n})\otimes M_{m}+\tau I_{n}\otimes{\color[rgb]{0,0,0}K_{m}}&\\ &(\Lambda_{\epsilon,n}+\alpha I_{n})\otimes M_{m}+\tau I_{n}\otimes{\color[rgb]{0,0,0}K_{m}}\end{bmatrix}^{-1}\widehat{\mathbf{v}},$